OpenStudy (anonymous):

he volume of a prism is found by multiplying the area of the base by the height. A rectangular prism has a volume of 144cm^3, and a square base with a width of 3 cm. What is the height of the prism, in cm? A. 4

OpenStudy (mathstudent55):

Since the base is a square, the length and width are the same, s. The height is h. \(V = s^2h\) \(144 = 3^2h \) \(9h = 144\) \(h = 16\) Answer the height is 16 cm.